### v1

# Quantum Advantage in Private Multiple Hypothesis Testing

Abstract
For multiple hypothesis testing based on classical data samples, we demonstrate a quantum advantage in the optimal privacy-utility trade-off (PUT), where the privacy and utility measures are set to (quantum) local differential privacy and the pairwise-minimum Chernoff information, respectively. To show the quantum advantage, we consider some class of hypotheses that we coin smoothed point masses. For such hypotheses, we derive an upper bound of the optimal PUT achieved by classical mechanisms, which is tight for some cases, and propose a certain quantum mechanism which achieves a better PUT than the upper bound. The proposed quantum mechanism consists of a classical-quantum channel whose outputs are pure states corresponding to a symmetric informationally complete positive operator-valued measure (SIC-POVM), and a depolarizing channel.
I Introduction
Statistical inference is essentially a process of collecting a large amount of data and extracting meaningful information from the collected data. If data is related to personal information, it is well known that there is a risk of other private information being leaked through this data [1, 2, 3, 4, 5]. For classical system, the amount of such a leakage can be bounded by perturbing the data with a privacy mechanism satisfying some privacy constraint. One of the most representative constraints is the local differential privacy (LDP) constraint [6, 7, 8], which is a slight variant of the differential privacy (DP) proposed by Dwork [9]. As proved by Issa et al. [10, Thm. 14], the LDP captures the limit of the probability of adversarial guessing on any kind of private information. However, in the sense of the data-processing inequality (DPI), the use of a privacy mechanism degrades the accuracy of inference, or the utility. Thus, one of the fundamental interests in private statistical inference is characterizing the optimal trade-off between the amount of leakage of private information and the utility, which is called the optimal privacy-utility trade-off (PUT). For classical system, the optimal PUTs for various inference tasks have been characterized either approximately or exactly, for example, [11] for discrete distribution estimation, [12] for mean estimation, and [13] for binary hypothesis testing.
On another note, in recent decades, quantum information processing has been receiving much attention as it can outperform classical information processing, i.e., there are quantum advantages in some tasks. For example, Grover’s algorithm [14] and Shor’s algorithm [15] have much lower algorithmic complexities than those of classical algorithms for searching and prime factorization, respectively. The BB84 protocol ensures much stronger security than classical protocols for sharing a secret key [16], and classical communication capacities of some multiple access channels can be enlarged [17]. On the other hand, it is also known that there is no quantum advantage for some tasks. For example, quantum cooperation has no advantage for some distributed compression [18] and nonlocal games [19], and algorithmic complexities of quantum machine learning models are comparable to those of classical machine learning models for some prediction tasks [20]. Proving the existence of a quantum advantage is valuable in the sense that we do not need to implement a quantum processing if there is no quantum advantage and an optimal classical processing is available. This perspective raises the question of whether there exists a quantum advantage in private statistical inference with respect to the optimal PUT.
In the past decade, the concept of LDP has been extended to quantum systems, which we call quantum LDP (QLDP) [21, 22, 23, 24]. The most relevant works with our work were done by Yoshida and Hayashi [22, 25]: Yoshida and Hayashi [22] showed that there is no quantum advantage in private statistical inference if original classical data is binary and the utility measure satisfies the DPI. This impossibility result was proved by showing that every $\epsilon$-QLDP mechanism is essentially classical, i.e., it is a composition of an $\epsilon$-LDP mechanism and a classical-quantum (CQ) channel, if classical data is binary ($\epsilon$ is a positive real number representing the limit of acceptable privacy leakage). Recently, Yoshida [25] showed that there is a quantum advantage in a certain optimization problem under privacy constraint.
In this paper, we demonstrate a quantum advantage in a certain operational scenario, a private multiple hypothesis testing where the original data is classical. The privacy constraint is set to QLDP, and the utility is chosen as the pairwise-minimum Chernoff information because it characterizes the rate of the optimal average error probability of testing as proved by Li [26]. Concerning some class of hypotheses that we call smoothed point masses, we demonstrate that the quantum advantage with respect to the optimal PUT exists. This result is proved by deriving an upper bound of the optimal classical PUT that is tight for some cases, and proposing a certain QLDP mechanism which achieves a better PUT than such an upper bound. To derive the upper bound of the optimal classical PUT, we exploit extremal mechanisms proposed by Kairouz et al. [13], and block design mechanisms proposed by Park et al. [27] shows the tightness of the bound for some cases. The proposed QLDP mechanism is a composition of a CQ channel whose outputs are pure states, and a depolarizing channel. Here, pure states and the level of depolarizing noise are suitably chosen to maximize the utility at some high privacy regime and for some hypotheses, while satisfying the QLDP constraint. The chosen pure states are states corresponding to a symmetric informationally complete positive operator-valued measure (SIC-POVM) [28, 29].
Notations: Throughout, let $\mathcal{H}_{d}$ denote a $d$-dimensional Hilbert space. The set of quantum states on a Hilbert space $\mathcal{H}$ is denoted by $\mathcal{D}(\mathcal{H})$. A measurement is described by a positive operator-valued measure (POVM) $\{\Lambda_{y}\}_{y\in\mathcal{Y}}$ where $\Lambda_{y}\geq 0$ and $\sum_{y\in\mathcal{Y}}\Lambda_{y}=I$. In addition, for a vector $q$ of length $d$, we write by $\mathrm{diag}(q)$ for a $d\times d$ diagonal matrix where elements $q_{x}$ for $x=1,\cdots,d$ correspond to diagonal elements. For $v\in\mathbb{N}$, we write by $[v]:=\{1,\ldots,v\}$ and $[0:v]:=\{0\}\cup[v]$. Let $\mathbbm{1}$ denote a matrix where all entries are $1$’s.
II Problem formulation
We consider a private multiple hypothesis testing scenario that consists of $n$ clients and a single server. We assume that for all $i\in[n]$, the $i$-th client has its own classical categorical data $X_{i}\in[v]$ for some $v\geq 2$, and $X_{i}$’s are i.i.d. following a distribution $P$ in some class of hypotheses $\mathbf{H}=\{P^{1},\cdots,P^{H}\}$. We assume that a hypothesis $P^{h}=(P_{1}^{h},\ldots,P_{v}^{h})$ is true with probability $\gamma_{h}>0$ for all $h\in[H]$, and denote $\gamma=(\gamma_{1},\ldots,\gamma_{H})$.
Instead of revealing original data, each of the clients perturbs its data into a quantum state through a classical-quantum (CQ) channel $Q:[v]\rightarrow\mathcal{D}(\mathcal{H}_{d})$ for some $d\geq 2$, which we call a privacy mechanism, and reveals the resultant quantum state to the server. For simplicity, we denote $Q_{x}:=Q(x)$. In order to limit the amount of leakage of private information, we impose the quantum local differential privacy (QLDP) constraint [21, 22, 23] on $Q$.
Definition 1.
For $\epsilon>0$, a privacy mechanism $Q$ is called an $\epsilon$-QLDP mechanism if for all POVM $\Lambda=\{\Lambda_{y}\}_{y\in\mathcal{Y}}$ and $x,x^{\prime}\in[v],y\in\mathcal{Y}$,
$\mathrm{Tr}(\Lambda_{y}Q_{x})\leq e^{\epsilon}\mathrm{Tr}(\Lambda_{y}Q_{x^{\prime}}).$ (1)
Note that $Q$ is an $\epsilon$-QLDP mechanism if and only if $Q_{x}\leq e^{\epsilon}Q_{x^{\prime}}$ for all $x,x^{\prime}\in[v]$, and $\epsilon$ can be interpreted as the maximum amount of leakage of private information through $Q$. The server tries to infer which hypothesis $P^{h}\in\mathbf{H}$ is the true hypothesis based on the quantum states revealed by all $n$ clients. Note that if $P^{h}$ is the true hypothesis, then the revealed states can be described as $\rho_{h}^{\otimes n}$, where
$\rho_{h}:=\sum\limits_{x\in[v]}P^{h}_{x}Q_{x}.$ (2)
The server performs the POVM $\{\Lambda_{h}\}_{h\in[H]}$ on $\rho_{h}^{\otimes n}$ which minimizes the average error probability. The minimum average error probability can be written as
$P_{e}^{*}(Q;\mathbf{H},\gamma)=\inf\limits_{\{\Lambda_{h}\}_{h\in[H]}:\text{POVM}}\sum\limits_{h\neq h^{\prime}}\mathrm{Tr}(\gamma_{h}\rho_{h}^{\otimes n}\Lambda_{h^{\prime}}).$ (3)
The previous work by Li [26] showed that as $n\rightarrow\infty$, $P_{e}^{*}$ exponentially converges to zero with the rate of the pairwise-minimum Chernoff information, i.e.,
$\lim\limits_{n\rightarrow\infty}-\frac{1}{n}\log P_{e}^{*}(Q;\mathbf{H},\gamma)=\min\limits_{h\neq h^{\prime}}C(\rho_{h},\rho_{h^{\prime}}),$ (4)
where $C(\rho,\sigma)$ denotes the Chernoff information,
$C(\rho,\sigma):=-\log\min\limits_{s\in[0,1]}\mathrm{Tr}(\rho^{s}\sigma^{1-s}).$ (5)
Note that the rate does not depend on $\gamma$. Based on this fact, we set the utility $M(Q;\mathbf{H})$ of a privacy mechanism $Q$ for a given set of hypotheses $\mathbf{H}$ to the pairwise-minimum Chernoff information, i.e.,
$M(Q;\mathbf{H}):=\min\limits_{h\neq h^{\prime}}C(\rho_{h},\rho_{h^{\prime}}).$ (6)
By jointly considering the privacy and the utility, the optimal quantum privacy-utility trade-off (PUT) is defined as
$M_{\mathrm{Q}}(v,\epsilon;\mathbf{H}):=\sup\limits_{Q:\epsilon\text{-QLDP}}M(Q;\mathbf{H}).$ (7)
Note that the classical scenario is a special case of the quantum scenario described above: A classical privacy mechanism is a CQ channel whose outputs $\{Q_{x}\}_{x\in[v]}$ can be diagonalized simultaneously with respect to some orthonormal basis of $\mathcal{H}_{d}$, i.e., $Q_{x}=\mathrm{Diag}(q_{x})$ for some probability vector $q_{x}\in\mathbb{R}^{d}$. Then, a classical privacy mechanism $Q$ can be identified as the row stochastic matrix $q\in\mathbb{R}^{v\times d}$ where $q_{xy}$ is the $y$-th component of $q_{x}$. For a classical privacy mechanism, Definition 1 is reduced to the original definition of the local differential privacy (LDP) [6, 7, 8] given as follows.
Definition 2.
For $\epsilon>0$, a classical privacy mechanism $q\in\mathbb{R}^{v\times d}$ is an $\epsilon$-LDP mechanism if for all $x,x^{\prime}\in[v],y\in[d]$,
$q_{xy}\leq e^{\epsilon}q_{x^{\prime}y}.$ (8)
The utility (exponential rate of the minimum average error probability) of a classical privacy mechanism becomes a reduced version of (6) [30],
$M(q;\mathbf{H})=\min\limits_{h\neq h^{\prime}}C(\tilde{q}_{h},\tilde{q}_{h^{\prime}}),$ (9)
where $\tilde{q}_{h}=q^{\top}P^{h}$ and
$C(\tilde{q}_{h},\tilde{q}_{h^{\prime}})=-\log\min\limits_{s\in[0,1]}\sum\limits_{y\in[d]}\tilde{q}_{hy}^{s}\tilde{q}_{h^{\prime}y}^{1-s}.$ (10)
Thus, the optimal classical PUT is defined as
$M_{\mathrm{C}}(v,\epsilon;\mathbf{H}):=\sup\limits_{q:\epsilon\text{-LDP}}M(q;\mathbf{H}).$ (11)
Clearly, we have $M_{\mathrm{Q}}\geq M_{\mathrm{C}}$. We say that there is a quantum advantage if $M_{\mathrm{Q}}>M_{\mathrm{C}}$.
About hypotheses, we focus on the one-parametric class of hypotheses $\mathbf{H}^{\eta}:=\{P^{h,\eta}\}_{h\in[H]}$ for $\eta\in[0,1]$, such that $H=v$ and
$P^{h,\eta}_{x}:=\eta\delta_{h,x}+\frac{1-\eta}{v},$ (12)
which we call smoothed point mass. Accordingly, we simply write $M^{\eta}(Q):=M(Q;\mathbf{H}^{\eta})$, $M_{\mathrm{Q}}^{\eta}(v,\epsilon):=M_{\mathrm{Q}}^{\eta}(v,\epsilon;\mathbf{H}^{\eta})$, and $M_{\mathrm{C}}^{\eta}(v,\epsilon):=M_{\mathrm{C}}^{\eta}(v,\epsilon;\mathbf{H}^{\eta})$. The goal of this work is to show $M_{\mathrm{Q}}^{\eta}>M_{\mathrm{C}}^{\eta}$.
Remark 1.
For $v=2$, there is no quantum advantage, i.e., $M_{\mathrm{Q}}(2,\epsilon;\mathbf{H})=M_{\mathrm{C}}(2,\epsilon;\mathbf{H})$ for all hypotheses $\mathbf{H}$ and $\epsilon>0$. This fact is a direct consequence of the results by Yoshida and Hayashi [22]: every $\epsilon$-QLDP mechanism is a sequential composition of an $\epsilon$-LDP mechanism and a CQ channel, or simply, an essentially classical $\epsilon$-QLDP mechanism. Because the Chernoff information satisfies the DPI [31, Chap. 4.4.1], there is no quantum advantage if $v=2$.
Remark 2.
The comparison between $M_{\mathrm{Q}}$ and $M_{\mathrm{C}}$ is fair in the sense that an $\epsilon$-QLDP mechanism ensures exactly the same level of privacy protection to an $\epsilon$-LDP mechanism in an operational sense. Operationally, $\epsilon$ captures the limit of the probability of adversarial guessing on any discrete private data based on the output of an $\epsilon$-LDP mechanism [10, Thm. 14]. When a client uses an $\epsilon$-QLDP mechanism for perturbing data, then an adversary should perform a measurement on an output quantum state to infer private information. As a direct consequence of [10, Thm. 14], $\epsilon$ captures the limit of the probability of adversarial guessing on private data for all possible measurements. The precise statement is as follows.
Proposition 1.
Let $X,Z$ be discrete random variables, and $Y$ be the measurement outcome corresponding to a POVM $\{\Lambda_{y}\}_{y\in\mathcal{Y}}$. Then,
$\sup\limits_{\{\Lambda_{y}\}_{y\in\mathcal{Y}}:\textnormal{POVM}}\max\limits_{y,x,x^{\prime}}\log\frac{\mathrm{Tr}(\Lambda_{y}Q_{x})}{\mathrm{Tr}(\Lambda_{y}Q_{x^{\prime}})}=\sup\limits_{\{\Lambda_{y}\}_{y\in\mathcal{Y}}:\textnormal{POVM}}\sup\limits_{P_{\hat{Z}|Y},P_{X,Z}}\log\frac{\max\limits_{y}\mathrm{Pr}(Z=\hat{Z}|Y=y)}{\max\limits_{z}\mathrm{Pr}(Z=z)}.$ (13)
III Main Results
Our main contribution is in demonstrating a quantum advantage in a private multiple hypothesis test.
Theorem 2.
For all $3\leq v\leq 9$, there exists $\epsilon_{1}>0$ such that
$\forall\epsilon\in(0,\epsilon_{1}],\quad M^{1}_{\mathrm{Q}}(v,\epsilon)>M^{1}_{\mathrm{C}}(v,\epsilon).$ (14)
Moreover, if $v=4$ or $9$, then there exist $\epsilon_{0}>0$ and $\eta_{0}<1$ such that
$\forall\epsilon\in(0,\epsilon_{0}],\eta\in[\eta_{0},1],\quad M_{\mathrm{Q}}^{\eta}(v,\epsilon)>M_{\mathrm{C}}^{\eta}(v,\epsilon).$ (15)
We prove the theorem by deriving an upper bound $\bar{M}_{\mathrm{C}}^{\eta}$ of $M^{\eta}_{\mathrm{C}}$ which is tight at $\eta=1$, and proposing a certain $\epsilon$-QLDP mechanism $Q$ with $\lim_{\epsilon\rightarrow 0}M^{1}(Q)/M^{1}_{C}(v,\epsilon)>1$. If $v=4$ or $9$, then this inequality can be extended by continuity for all $\eta$ close to $1$. The proposed $\epsilon$-QLDP mechanism is a sequential composition of a CQ channel whose outputs are pure states, followed by a depolarizing channel. Here, the amount of depolarizing noise is optimized to achieve the best utility while ensuring that the proposed mechanism is an $\epsilon$-QLDP mechanism. Also, pure states are chosen as states corresponding to a SIC-POVM [28, 29], because they induce the best utility for $\epsilon\approx 0$ and $\eta=1$. More details are described in the following subsections.
Definition 3.
An SIC-POVM is a set of $d^{2}$ unit vectors {|$\{\ket{\psi_{x}}\}_{x\in[d^{2}]}$ in $\mathcal{H}_{d}$ satisfying
$\sum\limits_{x=1}^{d^{2}}\ket{{\psi_{x}}}\!\bra{{\psi_{x}}}=dI_{d},$ (16)
and
$\forall x\neq x^{\prime},\quad|\left\langle{\psi_{x}}\middle|{\psi_{x^{\prime}}}\right\rangle|^{2}=\frac{1}{d+1}.$ (17)
In Fig. 1, we numerically plot $M^{\eta}(Q)/\bar{M}_{\mathrm{C}}^{\eta}(v,\epsilon)$ for $v=4,9$, where $Q$ is our proposed $\epsilon$-QLDP mechanism. Though Theorem 2 does not tell how small or large the values of $\epsilon_{1},\epsilon_{0},\eta_{0}$ are, the numerical results show that they are not negligible: $\epsilon_{1},\epsilon_{0}\geq 1$ and $\eta_{0}\leq 0.91$.
Figure 1: Comparison between $M^{\eta}(Q)$ and $\bar{M}_{\mathrm{C}}^{\eta}(v,\epsilon)$ for the proposed $\epsilon$-QLDP mechanism $Q$.
III-A Optimal classical PUT
In this subsection, we consider a classical scenario only and derive an upper bound of $M_{\mathrm{C}}^{\eta}$ which is tight at $\eta=1$. A high privacy regime $\epsilon\approx 0$ is of special interest, and for simple notation, we write $f(\epsilon)\overset{\epsilon\rightarrow 0}{\approx}g(\epsilon)$ if $\lim_{\epsilon\rightarrow 0}f(\epsilon)/g(\epsilon)=1$.
Proposition 3.
For all $v\geq 2,\epsilon>0$, and $\eta\in[0,1]$ we have
$M_{\mathrm{C}}^{\eta}(v,\epsilon)\leq\bar{M}_{\mathrm{C}}^{\eta}(v,\epsilon):=-\log\left(1-\frac{(v+\eta^{2}-1)(e^{\epsilon/2}-1)^{2}}{v(v-1)}\max\limits_{k\in[0:v]}\frac{k(v-k)}{ke^{\epsilon}+v-k}\right).$ (18)
Moreover, the equality holds if $\eta=1$ and
$M_{\mathrm{C}}^{1}(v,\epsilon)\overset{\epsilon\rightarrow 0}{\approx}\begin{cases}\frac{v}{16(v-1)}\epsilon^{2}&\textnormal{if }v=\textnormal{even}\\ \frac{v+1}{16v}\epsilon^{2}&\textnormal{if }v=\textnormal{odd}\end{cases}.$ (19)
The proof of the proposition consists of the converse part and the achievability part. Concerning the converse part, we show that it is enough to consider extremal mechanisms [13], and upper-bound the utility of an extremal mechanism using some techniques used in the proof of [25, Lem. 5.3]. For the achievability part, the utility of a block design mechanism [27] is shown to meet the converse bound at $\eta=1$. The proof of (19) is based on Taylor’s approximation.
III-A1 Converse
The concept of an extremal mechanism was originally proposed by Kairouz et al. [13]. The authors showed that if the utility satisfies some specific conditions, then it is enough to consider only extremal mechanisms for maximizing the utility over all LDP mechanisms. We extend the previous result to more general utilities, that satisfy the DPI. Formal descriptions are as follows.
Definition 4.
A real-valued function $U$ on the set of all row stochastic matrices $q$ is said to satisfy the DPI if $U(q)\geq U(q\Phi)$ for all row stochastic matrices $\Phi$.
Definition 5.
For an integer $v\geq 2$, let
$S^{(v)}:=\mathbbm{1}_{v\times 2^{v}}+(e^{\epsilon}-1)\mathrm{BIN}^{(v)},$ (20)
and $\mathrm{BIN}^{(v)}\in\{0,1\}^{v\times 2^{v}}$ be the matrix whose $y$-th column is the binary representation of $y-1$. A row stochastic matrix $q\in\mathbb{R}^{v\times{2^{v}}}$ is an extremal $\epsilon$-LDP mechanism if there exists a non-negative vector $\theta\in\mathbb{R}^{2^{v}}$ such that $q=S^{(v)}\mathrm{diag}(\theta)$ and $S^{(v)}\theta=\mathbbm{1}_{v}$.
Clearly, an extremal $\epsilon$-LDP mechanism is in fact an $\epsilon$-LDP mechanism. The proof of the following lemma is in Appendix.
Lemma 4.
Suppose a real-valued function $U$ on the set of all row stochastic matrices satisfies the DPI. Then, for all $v\geq 2$ and $\epsilon>0$, there exists an extremal $\epsilon$-LDP mechanism $q^{*}$ such that
$U(q^{*})=\sup\limits_{q:\epsilon\textnormal{-LDP}}U(q).$ (21)
Now, we prove the converse part of Proposition 3.
Proof:
Note that for probability vectors $(p,p^{\prime})$, $\sum\limits_{y}p_{y}^{s}(p^{\prime})_{y}^{1-s}$ is jointly concave in $(p,p^{\prime})$ for $s\in[0,1]$ [31, Prop. 4.8]. This implies that Chernoff information satisfies the DPI, and so does $M^{\eta}$. Therefore, Lemma 4 tells that $M_{\mathrm{C}}^{\eta}(v,\epsilon)=M(q)$ for some extremal $\epsilon$-LDP mechanism $q=S^{(v)}\mathrm{diag}(\theta)$. Thus,
$M_{\mathrm{C}}^{\eta}(v,\epsilon)$ $=M(q;\mathbf{H}^{\eta})=-\log\left(\max\limits_{h\neq h^{\prime}}\min\limits_{s\in[0,1]}\sum\limits_{y}\tilde{q}_{hy}^{s}\tilde{q}_{h^{\prime}y}^{1-s}\right)$ (22)
$\leq-\log\left(\max\limits_{h\neq h^{\prime}}\min\limits_{s\in[0,1]}\sum\limits_{x,x^{\prime}}P^{h,\eta}_{x}P^{h^{\prime},\eta}_{x^{\prime}}\sum\limits_{y}q_{xy}^{s}q_{x^{\prime}y}^{1-s}\right)$ (23)
$\leq-\log\left(\max\limits_{h\neq h^{\prime}}\sum\limits_{x,x^{\prime}}P^{h,\eta}_{x}P^{h^{\prime},\eta}_{x^{\prime}}\min\limits_{s\in[0,1]}\sum\limits_{y}q_{xy}^{s}q_{x^{\prime}y}^{1-s}\right)$ (24)
$=-\log\left(\max\limits_{h\neq h^{\prime}}\sum\limits_{x,x^{\prime}}P^{h,\eta}_{x}P^{h^{\prime},\eta}_{x^{\prime}}\min\limits_{s\in[0,1]}\sum\limits_{y}\theta_{y}\left(S_{xy}^{(v)}\right)^{s}\left(S_{x^{\prime}y}^{(v)}\right)^{1-s}\right),$ (25)
where the first inequality follows from the joint concavity and Jensen’s inequality.
We first solve the minimization over $s$. Note that
$\chi_{h,h^{\prime},y}(s)$ $:=\sum\limits_{y}\theta_{y}\left(S^{(v)}_{hy}\right)^{s}\left(S^{(v)}_{h^{\prime}y}\right)^{1-s}$ (26)
is convex in $s$. Thus, if ${\partialderivative{\chi_{h,h^{\prime},y}}{s}}(1/2)=0$, then $\min_{s\in[0,1]}\chi_{h,h^{\prime},y}(s)=\chi_{h,h^{\prime},y}(1/2)$. We have
${\partialderivative{\chi_{h,h^{\prime},y}}{s}}(1/2)$ $=\sum\limits_{y}\theta_{y}\sqrt{\left(S^{(v)}_{hy}\right)\left(S^{(v)}_{h^{\prime}y}\right)}\log\frac{S^{(v)}_{hy}}{S^{(v)}_{h^{\prime}y}}$ (27)
$=\epsilon e^{\epsilon/2}\left(\sum\limits_{y\in A^{e^{\epsilon},1}_{h,h^{\prime}}}\theta_{y}-\sum\limits_{y\in A^{1,e^{\epsilon}}_{h,h^{\prime}}}\theta_{y}\right),$ (28)
where
$A_{h,h^{\prime}}^{c,c^{\prime}}:=\left\{y:S^{(v)}_{hy}=c,S^{(v)}_{h^{\prime}y}=c^{\prime}\right\}.$ (29)
The following lemma proves ${\partialderivative{\chi_{h,h^{\prime},y}}{s}}(1/2)=0$, whose proof is in Appendix.
Lemma 5.
Suppose $S^{(v)}\theta=\mathbbm{1}_{v}$. Then, for all $h\neq h^{\prime}$ and $\epsilon>0$, we have
$\sum\limits_{y\in A^{e^{\epsilon},1}_{h,h^{\prime}}}\theta_{y}=\sum\limits_{y\in A^{1,e^{\epsilon}}_{h,h^{\prime}}}\theta_{y}.$ (30)
Now, we use similar techniques to those in the proof of [25, Lem. 5.3]. Based on the above argument, we get
$M_{\mathrm{C}}^{\eta}(v,\epsilon)$ $\leq-\log\left(\max\limits_{h\neq h^{\prime}}\sum\limits_{x,x^{\prime}}P^{h,\eta}_{x}P^{h^{\prime},\eta}_{x^{\prime}}\sum\limits_{y}\theta_{y}\sqrt{S_{xy}^{(v)}S_{x^{\prime}y}^{(v)}}\right)$ (31)
$\leq-\log\left(\frac{1}{v(v-1)}\sum\limits_{h\neq h^{\prime}}\sum\limits_{x,x^{\prime}}P^{h,\eta}_{x}P^{h^{\prime},\eta}_{x^{\prime}}\sum\limits_{y}\theta_{y}\sqrt{S_{xy}^{(v)}S_{x^{\prime}y}^{(v)}}\right)$ (32)
$=-\log\left(\frac{1}{v(v-1)}\left(\sum\limits_{h\neq h^{\prime}}\sum\limits_{x}P^{h,\eta}_{x}P^{h^{\prime},\eta}_{x}\sum\limits_{y}\theta_{y}S_{xy}^{(v)}+\sum\limits_{y}\theta_{y}\sum\limits_{x\neq x^{\prime}}\sum\limits_{h\neq h^{\prime}}P^{h,\eta}_{x}P^{h^{\prime},\eta}_{x^{\prime}}\sqrt{S_{xy}^{(v)}S_{x^{\prime}y}^{(v)}}\right)\right)$ (33)
$=-\log\left(\frac{1}{v(v-1)}\left(\sum\limits_{h\neq h^{\prime}}\langle P^{h,\eta},P^{h^{\prime},\eta}\rangle+\sum\limits_{y}\theta_{y}\sum\limits_{x\neq x^{\prime}}\sum\limits_{h\neq h^{\prime}}P^{h,\eta}_{x}P^{h^{\prime},\eta}_{x^{\prime}}\sqrt{S_{xy}^{(v)}S_{x^{\prime}y}^{(v)}}\right)\right),$ (34)
where the last equality follows from $\sum_{y}\theta_{y}S_{xy}^{(v)}=\sum_{y}q_{xy}=1$. By the definition of $P^{h,\eta}$, we have
$P^{h,\eta}_{x}P^{h^{\prime},\eta}_{x^{\prime}}=\eta^{2}\delta_{h,x}\delta_{h^{\prime},x^{\prime}}+\frac{1-\eta}{v}(\delta_{h,x}+\delta_{h^{\prime},x^{\prime}})+\frac{(1-\eta)^{2}}{v^{2}},$ (35)
which implies
$\sum\limits_{h\neq h^{\prime}}P^{h,\eta}_{x}P^{h^{\prime},\eta}_{x^{\prime}}=\frac{v+\eta^{2}-1}{v},$ (36)
for $x\neq x^{\prime}$, and
$\sum\limits_{h\neq h^{\prime}}\langle P^{h,\eta},P^{h^{\prime},\eta}\rangle=(1-\eta^{2})(v-1).$ (37)
Hence,
$M_{\mathrm{C}}^{\eta}(v,\epsilon)$ $\leq-\log\left(\frac{1-\eta^{2}}{v}+\frac{v+\eta^{2}-1}{v^{2}(v-1)}\sum\limits_{y}\theta_{y}\sum\limits_{x\neq x^{\prime}}\sqrt{S_{xy}^{(v)}S_{x^{\prime}y}^{(v)}}\right).$ (38)
Define $\Omega(y):=\left\{x:S^{(v)}_{xy}=e^{\epsilon}\right\}$ and $\beta_{k}:=\sum\limits_{y:|\Omega(y)|=k}\theta_{y}$. Then,
$\sum\limits_{y}\theta_{y}\sum\limits_{x\neq x^{\prime}}\sqrt{S_{xy}^{(v)}S_{x^{\prime}y}^{(v)}}$ $=\sum\limits_{k=0}^{v}\sum\limits_{y:|\Omega(y)|=k}\theta_{y}\sum\limits_{x\neq x^{\prime}}\sqrt{S^{(v)}_{xy}S^{(v)}_{x^{\prime}y}}$ (39)
$=\sum\limits_{k=0}^{v}\beta_{k}(k(k-1)e^{\epsilon}+2k(v-k)e^{\epsilon/2}+(v-k)(v-k-1)).$ (40)
Thus, we have
$M^{\eta}_{\mathrm{C}}(v,\epsilon)\leq-\log\left(\frac{1-\eta^{2}}{v}+\frac{v+\eta^{2}-1}{v^{2}(v-1)}\sum\limits_{k=0}^{v}\beta_{k}(k(k-1)e^{\epsilon}+2k(v-k)e^{\epsilon/2}+(v-k)(v-k-1))\right).$ (41)
Note that the constraint $S^{(v)}\theta=\mathbbm{1}_{v}$ implies $\sum\limits_{k=0}^{v}(ke^{\epsilon}+v-k)\beta_{k}=v$. Accordingly, $w=(w_{0},\ldots,w_{k})$, where $w_{k}:=(ke^{\epsilon}+v-k)\beta_{k}/v$, is a probability vector. Thus, we get
$M^{\eta}_{\mathrm{C}}(v,\epsilon)\leq-\log\left(\frac{1-\eta^{2}}{v}+\frac{v+\eta^{2}-1}{v(v-1)}\max\limits_{k\in[0:v]}\frac{k(k-1)e^{\epsilon}+2k(v-k)e^{\epsilon/2}+(v-k)(v-k-1)}{ke^{\epsilon}+v-k}\right),$ (42)
by maximizing (41) over all possible probability vectors $w$. Finally, a simple manipulation gives (18).
III-A2 Achievability
For an achievability part, we show that the utility of a block design mechanism [27] meets the upper bound derived in the converse part at $\eta=1$. A block design mechanism is constructed from a balanced incomplete block design (BIBD) [32, 33]. The symmetries of a BIBD (and a corresponding block design mechanism) meet the equality conditions of the inequalities in the proof of the converse part.
Definition 6.
A BIBD is a pair $(V,B)$ of a finite set of points $V$ and a non-empty set of blocks $B\subset 2^{V}$ satisfying the following symmetries for some non-negative integers $r,k,\lambda$:
  1. 1.
    $r$-regular: For all $p\in V$, |$|\{e\in B:p\in e\}|=r$,
  2. 2.
    $k$-uniform: For all $e\in B$, $|e|=k$,
  3. 3.
    $\lambda$-pairwise balanced: For all two distinct vertices $p,p^{\prime}$, we have |$|\{e\in B:\{p,p^{\prime}\}\subset e\}|=\lambda$.
We denote $v:=|V|$ and $b:=|B|$.
An $r$-regular, $k$-uniform, $\lambda$-pairwise balanced BIBD is conventionally written as $(v,k,\lambda)$-design, because $r,b$ can be determined by $(v,k,\lambda)$. Whenever $k\notin\{0,1,v\}$, the following identities should hold [32, Thm. 1.8, 1.9]:
$r=\frac{\lambda(v-1)}{k-1},\quad b=\frac{vr}{k}.$ (43)
If $k=0,1$, and $v$, then we must have $(r,b)=(0,1),(1,v)$, and $(1,1)$, respectively. Also, for any $k\in\{0,1,\cdots,v\}$, there exists a $(v,k,\lambda)$-design for some $\lambda$ (the complete $k$-uniform hypergraph).
Definition 7.
Let $(V,B)=([v],\{e_{1},\ldots,e_{b}\})$ be a $(v,k,\lambda)$-block design. A $(v,k,\lambda,\epsilon)$-block design mechanism constructed from $(V,B)$ is a row stochastic matrix $q\in\mathbb{R}^{v\times b}$ such that
$q_{xy}=\begin{cases}\frac{e^{\epsilon}}{re^{\epsilon}+b-r}&\textnormal{if }x\in e_{y}\\ \frac{1}{re^{\epsilon}+b-r}&\textnormal{if }x\not\in e_{y}\end{cases}.$ (44)
Clearly, $(v,k,\lambda,\epsilon)$-block design mechanism is an $\epsilon$-LDP mechanism. Note that a $(v,k,\lambda,\epsilon)$-block design mechanism constructed from the complete $k$-uniform hypergraph is known as the subset selection mechanism [11].
We conclude the subsection with the proof of the achievability part of Proposition 3 and (19).
Proof:
Let $q$ be a $(v,k,\lambda,\epsilon)$-block design mechanism. With the hypotheses $\mathbf{H}^{1}$, we have
$\tilde{q}_{hy}=q_{hy}=\begin{cases}\frac{e^{\epsilon}}{re^{\epsilon}+b-r}&\textnormal{if }h\in e_{y}\\ \frac{1}{re^{\epsilon}+b-r}&\textnormal{if }h\not\in e_{y}\end{cases}.$ (45)
Then, for all $h\neq h^{\prime}$, we have
$C(\tilde{q}_{h},\tilde{q}_{h^{\prime}})$ $=-\log\left(\min\limits_{s\in[0,1]}\sum\limits_{y}\tilde{q}_{hy}^{s}\tilde{q}_{h^{\prime}y}^{1-s}\right)$ (46)
$=-\log\left(\min\limits_{s\in[0,1]}\frac{\lambda e^{\epsilon}+(r-\lambda)\left(e^{\epsilon s}+e^{\epsilon(1-s)}\right)+b-2r+\lambda}{re^{\epsilon}+b-r}\right)$ (47)
$=-\log\left(\frac{\lambda e^{\epsilon}+2(r-\lambda)e^{\epsilon/2}+b-2r+\lambda}{re^{\epsilon}+b-r}\right),$ (48)
where the last equality is from the fact that $s=1/2$ is the solution of the minimization over $s\in[0,1]$ by the arithmetic mean-geometric mean (AM-GM) inequality. Combining with identities in (43), we get
$C(\tilde{q}_{h},\tilde{q}_{h^{\prime}})$ $=-\log\left(\frac{k(k-1)e^{\epsilon}++2k(v-k)e^{\epsilon/2}+(v-k)(v-k-1)}{(v-1)(ke^{\epsilon}+v-k)}\right)$ (49)
$=-\log\left(1-\frac{(e^{\epsilon/2}-1)^{2}}{v-1}\cdot\frac{k(v-k)}{ke^{\epsilon}+v-k}\right).$ (50)
By choosing $k\in[0:v]$ which maximizes the above, we get the desired lower bound of $M^{1}$ which meets the upper bound (18) at $\eta=1$.
For proving (19), let
$f(k)=\frac{k(v-k)}{ke^{\epsilon}+v-k},$ (51)
and $k^{*}=\operatorname*{arg\,max}\limits_{k\in[0:v]}f(k)$. By taking derivative in $k$,
$f^{\prime}(k)=\frac{e^{\epsilon}k^{2}-(v-k)^{2}}{(ke^{\epsilon}+v-k)^{2}}.$ (52)
Because $f^{\prime}(0)<0$, $f^{\prime}(v)>0$, and $f^{\prime}(k)=0$ has a unique solution $v/(e^{\epsilon/2}+1)$ in $k\in[0,v]$, we have
$k^{*}\in\left\{\left\lfloor\frac{v}{e^{\epsilon/2}+1}\right\rfloor,\left\lceil\frac{v}{e^{\epsilon/2}+1}\right\rceil\right\}.$ (53)
For $\epsilon\approx 0$, $k^{*}\in\{(v-1)/2,(v+1)/2\}$ if $v$ is odd, and $k^{*}\in\{v/2,v/2-1\}$ if $v$ is even. By comparing the values of $f$ at the candidate values of $k^{*}$, we have that if $\epsilon\approx 0$,
$k^{*}=\begin{cases}v/2&\text{if }v=\text{even}\\ (v-1)/2&\text{if }v=\text{odd}\end{cases}.$ (54)
Thus, Talyor’s approximation gives
$M_{C}^{1}(v,\epsilon)\overset{\epsilon\rightarrow 0}{\approx}\frac{k^{*}(v-k^{*})}{4v(v-1)}\epsilon^{2}=\begin{cases}\frac{v}{16(v-1)}\epsilon^{2}&\text{if }v=\text{even}\\ \frac{v+1}{16v}\epsilon^{2}&\text{if }v=\text{odd}\end{cases}.$ (55)
III-B Proposed QLDP mechanism
In this subsection, we consider a class of QLDP mechanisms which are sequential compositions of a CQ channel whose outputs are pure states, and a depolarizing channel. Precisely, we consider a privacy mechanism $Q$ in the following form:
$\forall x\in[v],\quad Q_{x}=\frac{\mu}{d}I_{d}+(1-\mu)\ket{{\psi_{x}}}\!\bra{{\psi_{x}}},$ (56)
where each of |$\ket{\psi_{x}}$ is a unit vector in $\mathcal{H}_{d}$ and $\mu\in[0,\frac{d}{d-1}]$. [^1] The depolarizing channel with parameter $\mu$ is a valid quantum channel (trace-preserving and completely positive) if and only if $\mu\in[0,\frac{d^{2}}{d^{2}-1}]$. Though, we allow $\frac{d^{2}}{d^{2}-1}<\mu\leq\frac{d}{d-1}$ since $Q$ a valid classical-quantum channel if and only if $\mu\in[0,\frac{d}{d-1}]$. It can be easily observed that $Q_{x}$ is positive semidefinite and $\mathrm{Tr}(Q_{x})=1$. This class of privacy mechanisms with $\mu\in(0,1]$ was previously considered by Yoshida [25]. We further choose suitable depolarizing parameter $\mu$ and pure states {|$\{\ket{\psi_{x}}\}_{x\in[v]}$ which maximize the utility at $\eta=1$ and $\epsilon\approx 0$, while ensuring the $\epsilon$-QLDP constraint. The chosen privacy mechanism shows a quantum advantage at $13\leq v\leq 9,\eta=$ 1, and $\epsilon\approx 0$. If $v=4$ or $9$, then the advantage can be extended to all $\eta$ close to $1$ by the continuity of $M^{\eta}$ and $\bar{M}_{\mathrm{C}}^{\eta}$.
We first derive a sufficient and necessary condition for such a privacy mechanism to be an $\epsilon$-QLDP mechanism, whose proof is similar to that of [25, Lem. 5.5].
Proposition 6.
For any given unit vectors {|$\{\ket{\psi_{x}}\}_{x\in[v]}\subset\mathcal{H}_{d}$, a privacy mechanism $Q$ in the form of (56) is an $\epsilon$-QLDP mechanism if and only if
$\mu\in\left[\frac{dg^{-}_{*}}{dg^{-}_{*}-1},\frac{dg^{+}_{*}}{dg^{+}_{*}-1}\right],$ (57)
where
$c_{x,x^{\prime}}:=|\left\langle{\psi_{x}}\middle|{\psi_{x^{\prime}}}\right\rangle|^{2},\quad c_{*}:=\min_{x\neq x^{\prime}}c_{x,x^{\prime}},\quad g^{\pm}_{*}:=\frac{1\pm\sqrt{1+\frac{1-c_{*}}{\sinh^{2}(\epsilon/2)}}}{2}.$ (58)
Proof:
By definitions, $Q$ is an $\epsilon$-QLDP mechanism if and only if for all $x,x^{\prime}\in[v]$,
$e^{\epsilon}Q_{x}\geq Q_{x^{\prime}}$ $\Leftrightarrow\frac{\mu(e^{\epsilon}-1)}{d}I_{d}\geq(1-\mu)(\ket{{\psi_{x^{\prime}}}}\!\bra{{\psi_{x^{\prime}}}}-e^{\epsilon}\ket{{\psi_{x}}}\!\bra{{\psi_{x}}})$ (59)
$\Leftrightarrow I_{d}\geq\frac{d(\mu-1)}{\mu(e^{\epsilon}-1)}(e^{\epsilon}\ket{{\psi_{x}}}\!\bra{{\psi_{x}}}-\ket{{\psi_{x^{\prime}}}}\!\bra{{\psi_{x^{\prime}}}}).$ (60)
[…]
$T_{x,x^{\prime}}=\frac{1}{e^{\epsilon}-1}\begin{pmatrix}e^{\epsilon}-c_{x,x^{\prime}}&-\sqrt{c_{x,x^{\prime}}(1-c_{x,x^{\prime}})}\\ -\sqrt{c_{x,x^{\prime}}(1-c_{x,x^{\prime}})}&c_{x,x^{\prime}}-1\end{pmatrix},$ (61)
and all other entries are zeros. Thus, we have
$e^{\epsilon}Q_{x}\geq Q_{x^{\prime}}$ $\Leftrightarrow I_{2}\geq\frac{d(\mu-1)}{\mu}T_{x,x^{\prime}}.$ (62)
For any any $2\times 2$ matrix $T$, its eigenvalues are given by
$\frac{\mathrm{Tr}(T)\pm\sqrt{\mathrm{Tr}(T)^{2}-4\mathrm{Det}(T)}}{2}.$ (63)
Because $\mathrm{Tr}(T_{x,x^{\prime}})=1$ and $\mathrm{Det}(T_{x,x^{\prime}})=\frac{c_{x,x^{\prime}}-1}{4\sinh^{2}(\epsilon/2)}$, the eigenvalues of $T_{x,x^{\prime}}$ are
$g^{\pm}_{x,x^{\prime}}:=\frac{1\pm\sqrt{1+\frac{1-c_{x,x^{\prime}}}{\sinh^{2}(\epsilon/2)}}}{2}.$ (64)
Combining with (62), the privacy mechanism satisfies $\epsilon$-QLDP if and only if for all $x,x^{\prime}\in[v]$ and $g=g_{x,x^{\prime}}^{\pm}$,
$1\geq\frac{d(\mu-1)}{\mu}g.$ (65)
Because $g^{-}_{x,x^{\prime}}\leq 0$ and $g^{+}_{x,x^{\prime}}\geq 1$, the above condition is equivalent to
$\forall x,x^{\prime},\quad\mu\in\left[\frac{dg^{-}_{x,x^{\prime}}}{dg^{-}_{x,x^{\prime}}-1},\frac{dg^{+}_{x,x^{\prime}}}{dg^{+}_{x,x^{\prime}}-1}\right].$ (66)
The remained part concerning $g_{*}^{\pm}$ follows from the fact that $dg/(dg-1)$ is decreasing in $g$ for $g\leq 0$ and $g\geq 1$. ∎
Now, we find the best $\mu$ and {|$\{\ket{\psi_{x}}\}_{x\in[v]}$ which maximize the utility at $\epsilon\approx 0$ and $\eta=1$, while satisfying the QLDP constraint in Proposition 6. It can be easily guessed that for a given {|$\{\ket{\psi_{x}}\}_{x\in[v]}$, the minimum $\mu$ in (57) would give the best utility because it induces the minimum level of depolarizing noise. A non-trivial part is in choosing pure states {|$\{\ket{\psi_{x}}\}_{x\in[v]}$. The Chernoff information can be interpreted as a distance between the states, thus the utility might be maximized if the minimum distance between the pure states is maximized. One of the candidates that achieve max-min distance is the set of uniformly separated states. On the other hand, the smallest level of depolarizing noise is attained when the maximum distance (with respect to the cosine distance) between the states is minimized, as in Proposition 6. Clearly, identical states achieve the min-max distance. For choosing the best states, we need to analyze how much influence that minimum and maximum distances have on the utility. The following proposition shows that at $\eta=1$ and $\epsilon\approx 0$, uniformly separated states induce the best utility.
Proposition 7.
For any given {|$\{\ket{\psi_{x}}\}_{x\in[v]}\subset\mathcal{H}_{d}$, and a privacy mechanism $Q$ in the form of (56), we have that
$\max\limits_{\mu}M^{1}(Q)$ $=-\log\left[c^{*}+\frac{1-c^{*}}{d}\left((d-2)\mu_{*}+2\sqrt{\mu_{*}(d-(d-1)\mu_{*})}\right)\right]$ (67)
$\overset{\epsilon\rightarrow 0}{\approx}\frac{1-c^{*}}{4d(1-c_{*})}\epsilon^{2}\leq\frac{\epsilon^{2}}{4d},$ (68)
where the maximum over $\mu$ is on (57), $\mu_{*}:=\frac{dg_{*}^{-}}{dg_{*}^{-}-1}$, and $c^{*}:=\max_{x\neq x^{\prime}}c_{x,x^{\prime}}$. The last inequality becomes equality if and only if $c_{x,x^{\prime}}$ is constant.
Proof:
Because $\rho_{h}=Q_{h}$ for $\eta=1$,
$\max\limits_{\mu}M^{1}(Q)$ $=\max\limits_{\mu}\min\limits_{h\neq h^{\prime}}C(Q_{h},Q_{h^{\prime}})$ (69)
$=-\log\left(\min\limits_{\mu}\max\limits_{h\neq h^{\prime}}\min\limits_{s\in[0,1]}\mathrm{Tr}(Q_{h}^{s}Q_{h^{\prime}}^{1-s})\right).$ (70)
Considering the spectral decomposition of $Q_{h}$, we have that
$Q_{h}^{s}=\left(\frac{\mu}{d}\right)^{s}I_{d}+\left(\left(\frac{\mu}{d}+1-\mu\right)^{s}-\left(\frac{\mu}{d}\right)^{s}\right)\ket{{\psi_{h}}}\!\bra{{\psi_{h}}}.$ | . (71)
Thus,
$\min\limits_{s\in[0,1]}\mathrm{Tr}\left(Q_{h}^{s}Q_{h^{\prime}}^{1-s}\right)$ (72)
$=c_{h,h^{\prime}}+\frac{(d-2)(1-c_{h,h^{\prime}})}{d}\mu+(1-c_{h,h^{\prime}})\min\limits_{s\in[0,1]}\left[\!\left(\frac{\mu}{d}\right)^{s}\left(\frac{\mu}{d}+1-\mu\right)^{1-s}+\left(\frac{\mu}{d}\right)^{1-s}\left(\frac{\mu}{d}+1-\mu\right)^{s}\!\right]$ (73)
$=c_{h,h^{\prime}}+\frac{1-c_{h,h^{\prime}}}{d}\left((d-2)\mu+2\sqrt{\mu(d-(d-1)\mu)}\right),$ (74)
where the last equality follows from the AM-GM inequality. By calculating partial derivatives, it can be shown that
$\frac{(d-2)\mu+2\sqrt{\mu(d-(d-1)\mu)}}{d}$ (75)
is concave in $\mu$ and its maximum value $1$ is attained at $\mu=1$. Thus, the right-hand side of (74) is increasing in $c_{h,h^{\prime}}\in[0,1]$. Accordingly, we have
$\max\limits_{h\neq h^{\prime}}\min\limits_{s\in[0,1]}\mathrm{Tr}(Q_{h}^{s}Q_{h^{\prime}}^{1-s})=c^{*}+\frac{1-c^{*}}{d}\left((d-2)\mu+2\sqrt{\mu(d-(d-1)\mu)}\right).$ (76)
Next, again by the concavity of (75), the minimizer $\mu$ that minimizes the above is one of the endpoints of (57), $\mu_{*}$ or $\mu_{*}^{+}=\frac{dg^{+}_{*}}{dg^{+}_{*}-1}$. We prove that actually $\mu=\mu_{*}$ is a minimizer, by showing
$\frac{(d-2)\mu_{*}+2\sqrt{\mu_{*}(d-(d-1)\mu_{*})}}{d}\leq\frac{(d-2)\mu_{*}^{+}+2\sqrt{\mu_{*}^{+}(d-(d-1)\mu_{*}^{+})}}{d}.$ (77)
Substituting $\mu=\frac{dg}{dg-1}$ in (75) gives
$\frac{(d-2)\mu+2\sqrt{\mu(d-(d-1)\mu)}}{d}|_{\mu=\frac{dg}{dg-1}}=\frac{(d-2)g}{dg-1}+\frac{2}{|dg-1|}\sqrt{g(g-1)}.$ (78)
Also, we can write
$g_{*}^{-}=-\alpha,\quad g_{*}^{+}=1+\alpha,$ (79)
where
$\alpha=\frac{\sqrt{1+\frac{1-c_{x,x^{\prime}}}{\sinh^{2}(\epsilon/2)}}-1}{2}\geq 0.$ (80)
Hence, (77) is equivalent to
$\frac{(d-2)\alpha}{d\alpha+1}+\frac{2}{d\alpha+1}\sqrt{\alpha(\alpha+1)}$ $\leq\frac{(d-2)(1+\alpha)}{d\alpha+d-1}+\frac{2}{d\alpha+d-1}\sqrt{\alpha(\alpha+1)}.$ (81)
Multiplying both sides by $(d\alpha+1)(d\alpha+d-1)>0$, above is equivalent to
$(d-2)\alpha(d\alpha+d-1)+2(d\alpha+d-1)\sqrt{\alpha(\alpha+1)}$ $\leq(d-2)(1+\alpha)(d\alpha+1)+2(d\alpha+1)\sqrt{\alpha(\alpha+1)}$ (82)
$\Leftrightarrow 2(d-2)\sqrt{\alpha(\alpha+1)}$ $\leq(d-2)(2\alpha+1).$ (83)
Since $d\geq 2$ and $(2\alpha+1)^{2}-(2\sqrt{\alpha(\alpha+1)})^{2}=1>0$, we have $12\sqrt{\alpha(\alpha+1)}\leq 2\alpha+$ 1. These prove (83), and implies (77) and (67).
For the last part of the proof, Taylor’s approximation and $c_{*}\leq c^{*}$ gives (68). It is trivial that the inequality in (68) becomes equality if and only if $c_{x,x^{\prime}}$ is constant. ∎
According to Proposition 7, we choose a set of pure states as a subset of a SIC-POVM in Definition 3. Also, it can be seen that smaller $d$ gives a better utility. Concerning this point, the number of pure states {|$\{\ket{\psi_{x}}\}$ in $\mathcal{H}_{d}$ with constant $c_{x,x^{\prime}}$ is at most $d^{2}$ [29, Thm. 2.28], and SIC-POVM in Definition 3 is such a maximal set. The existence of SIC-POVM is known up to $d\leq 151$, and it has been conjectured that SIC-POVM exists for all $d$ [34, 29, 28]. Combining with these facts and Propositions 3 and 7, we can show quantum advantage at $\eta=1$ as in the following corollary whose proof is straightforward.
Corollary 8.
Let {|$\{\ket{\psi_{x}}\}_{x\in[v]}\subset\mathcal{H}_{d}$ be a subset of a SIC-POVM. For a privacy mechanism $Q$ in the form of (56) with $\mu=\mu_{*}$, we have
$\lim\limits_{\epsilon\rightarrow 0}\frac{M^{1}(Q)}{M_{\mathrm{C}}^{1}(v,\epsilon)}\geq\begin{cases}\frac{4(v-1)}{v\lceil\sqrt{v}\rceil}&\textnormal{if}v=\textnormal{even}\\ \frac{4v}{(v+1)\lceil\sqrt{v}\rceil}&\textnormal{if }v=\textnormal{odd}\end{cases}.$ (84)
Thus, if $3\leq v\leq 9$, there exists $\epsilon_{1}>0$ such that $M^{1}(Q)>M^{1}_{\mathrm{C}}(v,\epsilon)$ for all $\epsilon\in(0,\epsilon_{1}]$.
If $v=d^{2}$ for some natural number $d\geq 2$, then the utility of the proposed privacy mechanism can be calculated as in the following corollary. We omit the proof because it is almost identical to (71)-(76), combined with Definition 3.
Corollary 9.
Let $v=d^{2}$ and {|$\{\ket{\psi_{x}}\}_{x\in[v]}\subset\mathcal{H}_{d}$ be a SIC-POVM. For a privacy mechanism $Q$ in the form of (56) with $\mu=\mu_{*}$, we have
$M^{\eta}(Q)=-\log\left[\frac{1}{d+1}\left(1+(d-2)\mu_{\eta}+2\sqrt{\mu_{\eta}(d-(d-1)\mu_{\eta})}\right)\right],$ (85)
where $\mu_{\eta}:=1-\eta+\mu_{*}\eta$.
Because $M^{\eta}(Q)$ and $\bar{M}_{\mathrm{C}}^{\eta}$ are continuous in $\eta$ for $v=d^{2}$, we can extend quantum advantage for all $\eta$ close to $1$ by continuity. Therefore, Propositions 36, and Corollaries 89 prove Theorem 2.
IV Conclusion & Discussions
We showed a quantum advantage in private multiple hypothesis testing with respect to the optimal PUT. Here, we considered $\epsilon$-QLDP constraint as a privacy constraint and smoothed point mass hypotheses $\mathbf{H}^{\eta}$ on categorical data with $v$ possible values. For showing this result, we derived an upper bound of the optimal classical PUT, and proposed a QLDP mechanism which is a sequential composition of a CQ channel whose outputs are pure states and a depolarizing channel. The depolarization parameter and pure states were chosen to maximize the utility for some cases, and the chosen states are elements of a SIC-POVM.
A promising future work would be on showing a quantum advantage for all $v,\eta,\epsilon$, or for more general hypotheses, or for other private statistical inference tasks. Concerning the first two directions, a block design mechanism [27] generalized for a quantum system can be considered. In more detail, we can consider a CQ channel whose outputs are mixtures of pure states {|$\{\ket{\psi_{e}}\}$ where the coefficients are determined by a block design (cf. Definition 7) as follows:
$Q_{x}=\frac{\mu}{r\mu+(b-r)(1-\mu)}\sum\limits_{e\in B:x\in e}\ket{{\psi_{e}}}\!\bra{{\psi_{e}}}+\frac{1-\mu}{r\mu+(b-r)(1-\mu)}\sum\limits_{e\in B:x\not\in e}\ket{{\psi_{e}}}\!\bra{{\psi_{e}}},$ (86)
where $([v],B)$ is a $(v,k,\lambda)$-design. However, there are technical difficulties in calculating the utility and checking the QLDP constraint, because it is difficult to specify the spectral decomposition of a general mixture of pure states. If there are some choices of block designs and pure states {|$\{\ket{\psi_{e}}\}$ that allow us to bypass such difficulties, they might show quantum advantages for more general cases. For the last direction, widely used inference tasks such as binary hypothesis testing and composite hypothesis testing (where utilities are set to relative entropy by Stein’s lemma [35]), and parametric estimation would be considered.
Proof:
It is enough to show that for any $\epsilon$-LDP mechanism $q\in\mathbb{R}^{v\times b}$, there exists an extremal $\epsilon$-LDP mechanism $q^{*}\in\mathbb{R}^{v\times 2^{v}}$ and a row stochastic matrix $\Phi\in\mathbb{R}^{2^{v}\times b}$ such that $q=q^{*}\Phi$. For given $\epsilon$-LDP mechanism $q$, we have $q=\bar{q}\cdot\mathrm{diag}(\theta)$, where $\theta\in[0,1]^{v}$, $\theta_{y}=\min_{x\in[v]}q_{xy}$, and $\bar{q}\in[0,e^{\epsilon}]^{v\times b}$. Let $\bar{q}^{y}$ be the $y$-th column of $\bar{q}$. Then, $\bar{q}^{y}$ is a convex combination of the columns of $S^{(v)}$, i.e., there exists $\bar{\theta}^{y}\in\mathbb{[}0,1]^{2^{v}}$ such that $\bar{q}^{y}=S^{(v)}\bar{\theta}^{y}$ and $\sum_{x\in 2^{v}}\bar{\theta}^{y}_{x}=1$. Thus,
$q=S^{(v)}\begin{bmatrix}\bar{\theta}^{1},\cdots,\bar{\theta}^{b}\end{bmatrix}\mathrm{diag}(\theta)=S^{(v)}\mathrm{diag}(\zeta)\bar{\zeta},$ (87)
where $\zeta\in[0,1]^{2^{v}}$, $\zeta_{z}=\sum_{y\in[b]}\theta_{y}\bar{\theta}^{y}_{z}$, and $\bar{\zeta}_{zy}=\theta_{y}\bar{\theta}^{y}_{z}/\zeta_{z}$. Clearly, $\bar{\zeta}$ is a row stochastic matrix. In addition, $q^{*}=S^{(v)}\mathrm{diag}(\zeta)$ is an extremal $\epsilon$-LDP mechanism because
$S^{(v)}\zeta=S^{(v)}\sum\limits_{y\in[b]}\theta_{y}\bar{\theta}^{y}=\sum\limits_{y\in[b]}\theta_{y}S^{(v)}\bar{\theta}^{y}=\sum\limits_{y\in[b]}\theta_{y}\bar{q}^{y}=q\theta=\mathbbm{1}_{v}.$ (88)
Proof:
Let $A_{h,h^{\prime}}^{c,c^{\prime}}$ be the set defined in (29). Then,
11 11 $=\left(S^{(v)}\theta\right)_{h}=\sum\limits_{y\in A_{h,h^{\prime}}^{e^{\epsilon},e^{\epsilon}}}e^{\epsilon}\theta_{y}+\sum\limits_{y\in A_{h,h^{\prime}}^{e^{\epsilon},1}}e^{\epsilon}\theta_{y}+\sum\limits_{y\in A_{h,h^{\prime}}^{1,e^{\epsilon}}}\theta_{y}+\sum\limits_{y\in A_{h,h^{\prime}}^{1,1}}\theta_{y}$ (89)
$=\left(S^{(v)}\theta\right)_{h^{\prime}}=\sum\limits_{y\in A_{h,h^{\prime}}^{e^{\epsilon},e^{\epsilon}}}e^{\epsilon}\theta_{y}+\sum\limits_{y\in A_{h,h^{\prime}}^{e^{\epsilon},1}}\theta_{y}+\sum\limits_{y\in A_{h,h^{\prime}}^{1,e^{\epsilon}}}e^{\epsilon}\theta_{y}+\sum\limits_{y\in A_{h,h^{\prime}}^{1,1}}\theta_{y}.$ (90)
Therefore, we get the desired result because
$(e^{\epsilon}-1)\left(\sum\limits_{y\in A_{h,h^{\prime}}^{e^{\epsilon},1}}\theta_{y}-\sum\limits_{y\in A_{h,h^{\prime}}^{1,e^{\epsilon}}}\theta_{y}\right)=0.$ (91)